Mr. Sunshine's favourite toy is a hot red toy car. On one sunny day he measured the final velocity of his car to be 10.0 m/s (W). If the toy car accelerated at a constant rate of 2.0 m/s^2 [W] over 12 m [W], find the initial velocity of the car.


1
Expert's answer
2021-09-09T10:51:22-0400
"v^2=v_0^2+2ad,""v_0=\sqrt{v^2-2ad},""v_0=\sqrt{(10\ \dfrac{m}{s})^2-2\cdot2\ \dfrac{m}{s^2}\cdot12\ m}=7.2\ \dfrac{m}{s}."